# Arithmetic in Transformers Explained

### v9

@@ Appendix G Appendix: Complexity
Appendix H Appendix: Addition Hypothesis 1
Given the 2-layer attention pattern’s similarity to 1-layer attention pattern, and the above evidence, our first (incorrect) hypothesis was that the 2-layer algorithm:
  • Is based on the same SA, SC and SS operations as the 1-layer.
  • Uses the new early positions to (somehow) do the SS calculations with higher accuracy than the 1-layer model.
  • The long double staircase still finalizes each answer digit’s calculation.
  • The two attention nodes in the long double staircase positions do the SA and SC calculations and pull in SS information calculated in the early positions.
If this is correct then the 2-layer algorithm successfully completes these calculations:
  • A0 = A0.SA
  • A1 = A1.SA + A0.SC
  • A2 = A2.SA + (A1.SC or (A1.SS & A0.SC))
  • A3 = A3.SA + (A2.SC or (A2.SS & A1.SC) or (A2.SS & A1.SS & A0.SC))
  • A4 = A4.SA + (A3.SC or (A3.SS & A2.SC) or (A3.SS & A2.SS & A1.SC) or (A3.SS & A2.SS & A1.SS & A0.SC))
  • A5 = A4.SC or (A4.SS & A3.SC) or (A4.SS & A3.SS & A2.SC) or (A4.SS & A3.SS & A2.SS & A1.SC) or (A4.SS & A3.SS & A2.SS & A1.SS & A0.SC)
Our intuition is that there are not enough useful nodes in positions 8 to 11 to complete the A5 calculation this way. So we abandoned this hypothesis.
Appendix I Appendix: Addition Hypothesis 2
Our second (incorrect) hypothesis was that the 2-layer algorithm has a more compact data representation, so it can pack more calculations into each node, allowing it to accurately predict A5 in step 11.
We claimed the model stores the sum of each digit-pair as a single token in the range “0” to “18” (covering 0+0 to 9+9). We name this operator $A_{n}$.T, where T stands for “token addition”:
  • $A_{n}$.T = $D_{n}$+ $D^{\prime}_{n}$
The $A_{n}$.T operation does not understand mathematical addition. Tab. 10 shows how the model implements the T operator as a bigram mapping.
Table 10: Implementing the T operator as a bigram mapping from 2 input tokens to 1 result token.
$D_{n}$vs $D^{\prime}_{n}$ 0 1 4 5 8 9
0 0 1 4 5 8 9
1 1 2 5 6 9 10
4 4 5 8 9 12 13
5 5 6 9 10 13 14
8 8 9 12 13 16 17
9 9 10 13 14 17 18
$A_{n}$.T is a compact way to store data. Tab. 11 show how, if it needs to, the model can convert a $A_{n}$.T value into a one-digit-accuracy SA, SC or SS value.)
Table 11: Converting a $A_{n}$.T value into a SA , SC or SS value.
$A_{n}$.T $A_{n}$.SA $A_{n}$.SC $A_{n}$.SS
0 0 0 0
1 1 0 0
8 8 0 0
9 9 0 1
10 0 1 0
17 7 1 0
18 8 1 0
Our notation shorthand for one-digit-accuracy these “conversion” bigram mappings is:
  • $A_{n}$.SA = ($A_{n}$.T % 10) where % is the modulus operator
  • $A_{n}$.SC = ($A_{n}$.T // 10) where // is the integer division operator
  • $A_{n}$.SS= ($A_{n}$.T == 9) where == is the equality operator
The A0.T value is accurate. But the other $A_{n}$.T values are not accurate because each is constrained to information from just one digit. We define another more accurate operator $A_{n}$.T2 that has “two-digit accuracy”. $A_{n}$.T2 is the pair sum for the nth digit plus the carry bit (if any) from the n-1th digit T:
  • $A_{n}$.T2 = $A_{n}$.T + $A_{n-1}$.SC
$A_{n}$.T2 is more accurate than $A_{n}$.T. The $A_{n}$.T2 value is always in the range “0” to “19” (covering 0+0+0 to 9+9+CarryOne). Tab. 12 show how the model can implement the T2 operator as a mapping.
$A_{n}$.T $A_{n}$.T2 if $A_{n-1}$.SC==0 $A_{n}$.T2 if $A_{n-1}$.SC==1
0 0 1
1 1 2
9 9 10
10 10 11
17 17 18
18 18 19
Table 12: Calculating $A_{n}$.T2 from $A_{n}$.T and $A_{n-1}$.T
Following this pattern, we define operators $A_{n}$.T3, $A_{n}$.T4 and $A_{n}$.T5 with 3, 4 and 5 digit accuracy respectively:
  • $A_{n}$.T3 = $A_{n}$.T + ( $A_{n-1}$.T2 // 10 ) aka $A_{n}$.T + $A_{n-1}$.SC2
  • $A_{n}$.T4 = $A_{n}$.T + ( $A_{n-1}$.T3 // 10 ) aka $A_{n}$.T + $A_{n-1}$.SC3
  • $A_{n}$.T5 = $A_{n}$.T + ( $A_{n-1}$.T4 // 10 ) aka $A_{n}$.T + $A_{n-1}$.SC4
The value A4.T5 is accurate as it integrates SC and cascading SS data all the way back to and including A0.T. The values A1.T2, A2.T3, A3.T4 are also all accurate. If the model knows these values it can calculate answer digits accurately:
  • A1 = A1.T2 % 10
  • A2 = A2.T3 % 10
  • A3 = A3.T4 % 10
  • A4 = A4.T5 % 10
  • A5 = A4.T5 // 10
In this hypothesis, all the answer digits are accurately calculated using the nodes in positions 8 to 11. This hypothesis 2 is feasible, elegant and compact - reflecting the authors (human) values for good code.
Experimenation shows the model does not implement this hypothesis. It retains the long staircase SA calculations in positions 11 to 16. Why? Two reasons suggest themselves:
  • Hypothesis 2 is too compact. The model is not optimising for compactness. The long staircase is discovered early in training, and it works for simple questions. Once the overall algorithm gives low loss consistently it stops optimising.
  • Hypothesis 2 accurately predicts all answer digits in step 11 - reflecting the authors (human) values for good code. The model is not motivated to do this. It just needs to accurately predict A5 as 1 or 0 in step 11 and A4 in step 12 - nothing more.
We abandoned this hypothesis.
Appendix J Appendix: Addition Hypothesis 3
  • Ablating the nodes one by one shows which answer digit(s) are reliant on each node (Ref Table 1). Most interestingly, ablating P10.L0.H1 impacts the answer digits A5, A4, A3, A2 (but not A1 and A0). This node is used in the calculation of A5, A4, A3, A2 in prediction steps 11, 12, 13 and 14. These relationships are constraints that are all obeyed by the pseudo-code.
  • The pseudo-code has 4 instances where $A_{n}.ST$ is calculated using TriCase. PCA of the corresponding nodes (P8.L0.H1, P9.L0.H1, P11.L0.H2 and P14.L0.H1) shows tri-state output for the specified $D_{n}$. (see Figure 9).
  • Activation patching (aka interchange intervention) experiments at attention head level confirmed some aspects of the calculations (see § K for details.
  • The pseudo code includes calculations like A1.ST which it says is calculated in P9.L0.H1 and P9.L0.MLP. Ablation tells us both nodes are necessary. For the attention head we use the PCA results for insights. We didn’t implement a similar investigative tool for the MLP layer, so in the pseudo-code we attribute the calculation of A1.ST to both nodes.
  • For P10.L0.H1, the attention head PCA could represent either a bi-state or tri-state output (see Figure 10). The MLP layer at P10.L0.MLP could map the attention head output to either a bi-state or tri-state. We cannot see which. The pseudo-code shows a tri-state calculation at P10.L0.MLP, but with small alterations the pseudo-code would work with a bi-state output.
  • For P15.L0.H1 the attention head PCA could represent either a bi-state or tri-state output (see Figure 10). The pseudo-code shows a bi-state calculation A0.SC at P15.L0.H1, but with small alterations the pseudo-code would work with a tri-state output.
  • The calculation of A1.ST2 in P14.L0.H1 is a interesting case. The model needs A1.ST2 for A2 accuracy. The model could simply reuse the accurate A1.ST2 value calculated in P10. Activation patching shows that it does not. Instead the P14 attention heads calculate A1.ST1 from D1 and D’1 directly, and only relies on the P10.D1.ST2 value in the case where A1.ST2 != A1.ST. That is, the calculation is “use P14.A1.ST1 value else use A1.ST2 values”. This aligns with the model learning the P10.A1.ST calculation early in training (for 90% accuracy) and later learning that P10.A1.ST2 contains additional information it can use to get to >99.999% accuracy.
Figure 6: For a sample 5-digit 2-layer 3-head addition model, this map shows a compacted view of all useful token positions (horizontally) and all useful attention heads and MLP layers (vertically) used in predictions as green cells. Each cell shows the simpliest (lowest complexity) quanta S0, S1, etc impacted when we ablate each node. To answer S0 questions, only the S0 nodes are used. To answer S1 questions, S0 and S1 nodes are used, etc. The model only uses nodes in nine token positions.
Figure 7: This map shows the % of enriched questions that fail when we ablate each node in a 5-digit 2-layer 3-head addition model. The model only uses nodes in token positions P8 to P16 (i.e. tokens D’2 to A1). Lower percentages correspond to rarer edge cases. The grey space represents nodes that are not used by the model.
Figure 8: This map shows the input tokens each attention head attends to at each token position in a 5-digit 2-layer 3-head addition model. At token position P11 the model predicts the first answer digit A5. All digit pairs (e.g. D2 D’2) are attended to by P11.
Figure 9: For 5-digit addition, for these attention heads, for exactly 1 answer digit $A_{n}$each, PCA shows these interpretable results. The dot colours show the TriCase value of each question. The PCA data and TriCase quanta are both tri-state and strongly correlated.
Figure 10: For 5-digit addition, for these attention heads, for exactly 1 digit $A_{n}$each, PCA shows these interpretable results. The dot colours show the TriCase value of each question. The PCA and TriCase data are strongly correlated, but the PCA data has 2 states.
Appendix K Appendix: Addition Interchange Interventions
To test the hypothesis 3 mapping of the mathematical framework (casual abstraction) to the model attention heads, various “interchange interventions” (aka activation patching) experiments were performed on the model, where
  • A particular claim about an attention head has selected for testing.
  • The model predicted answers for sample test questions, and the attention head activations were recorded (stored).
  • The model then predicted answers for more questions, but this time we intervened during the prediction to override the selected attention head activations with the activations from the previous run.
Using this approach we obtained the findings in Tab. 13.
Nodes Claim: Attention head(s) perform … Finding: Attention head(s) perform a function that …
P8.L0.H1 and MLP A2.ST = TriCase( D2, D’2 ) impacting A4 and A5 accuracy Based on D2 and D’2. Triggers on a A2 carry value. Provides carry bit used in A5 and A4 calculation.
P9.L0.H1 and MLP A1.ST = TriCase( D1, D’1 ) impacting A5, A4 and A3 accuracy Based on D1 and D’1. Triggers on a A1 carry value. Provides carry bit used in A5, A4 and A3 calculation.
P10.L0.H1 and MLP A1.ST2 = TriAdd( A1.ST, TriCase(D0, D’0) ) impacting A5, A4, A3 and A2 accuracy Based on D0 and D’0. Triggers on a A0 carry value. Provides carry bit used in A5, A4, A3 and A2 calculation.
P11.L0.H1 and MLP A3.ST4 = TriAdd( TriCase( D3, D’3 ), TriAdd( A2.ST, A1.ST2)) impacting A5 accuracy Based on D3 and D’3. Triggers on a A3 carry value. Provides carry bit used in A5 calculations.
P11.L0.H2 and MLP A4.ST = TriCase(D4, D’4) impacting A5 accuracy A4 Based on D4 and D’4. Triggers on a A4 carry value. Provides carry bit used in A5 calculation.
P12.L0.H0+H2 and MLP A4.SA = (D4 + D’4) % 10 impacting A4 accuracy A4 Sums D4 and D’4. Impacts A4.
P13.L0.H0+H2 and MLP A3.SA = (D3 + D’3) % 10 impacting A3 accuracy Sums D3 and D’3. Impacts A3.
P14.L0.H0+H2 and MLP A2.SA = (D2 + D’2) % 10 impacting A2 accuracy Sums D2 and D’2. Impacts A2.
P14.L0.H1 and MLP (D1 + D’1) / 10 + P10.A1.ST2 info impacting A2 accuracy Calculates P10.A1.ST1 but add P10.A1.ST2 info when A1.ST != A1.ST2. Impacts A2
P15.L0.H0+H2 and MLP A1.SA = (D1 + D’1) % 10 impacting A1 accuracy Sums D1 and D’1. Impacts A1.
P15.L0.H1 and MLP A0.SC = (D0 + D’0) / 10 impacting A1 accuracy Triggers when D0 + D’0 >10. Impacts A1 digit by 1
P16.L0.H0+H2 and MLP A0.SA = (D0 + D’0) % 10 impacting A0 accuracy Sums D0 and D’0. Impacts A0.
Table 13: Interchange Interventions experiments used activation patching to test the claims addition hypothesis 3 made for each attention head in a sample mixed model. Experimental results are consistent with hypothesis 3 for all nodes.
+ve Sub Contains -ve Sub Contains Like
M0 MD N/A N/A S0
M1 MD,MB N1 ND,NB S1
M2 MD,MBx2 N2 ND,NBx2 S2
M3 MD,MBx3 N3 ND,NBx3 S3
M4 MD,MBx4 N4 ND,NBx4 S4
Table 14: We define “positive-answer subtraction” and “negative-answer subtraction” calculation complexity quanta that parallel the addition quanta.
Appendix L Appendix: N-Digit Addition
@@ Appendix I Appendix: Addition Hypothesis 2
Appendix M Appendix: Mixed Model Initialization
@@
Appendix N Appendix: N-Digit Subtraction

### v10

@@ Appendix I Appendix: Complexity
Figure 10: For a sample 5-digit 2-layer 3-head addition model, this map shows a compacted view of all useful token positions (horizontally) and all useful attention heads and MLP layers (vertically) used in predictions as green cells. Each cell shows the simplest (lowest complexity) quanta S0, S1, etc impacted when we ablate each node. To answer S0 questions, only the S0 nodes are used. To answer S1 questions, S0 and S1 nodes are used, etc. The model only uses nodes in ten token positions.
Figure 11: This map shows the % of enriched questions that fail when we ablate each node in a 5-digit 2-layer 3-head addition model. The model only uses nodes in token positions P6 to P17 (i.e. tokens D’4 to A1). Lower percentages correspond to rarer edge cases. The grey space represents nodes that are not used by the model.
Figure 12: This map shows the input tokens each attention head attends to at each token position in a 5-digit 2-layer 3-head addition model. At token position P12 the model predicts the first answer digit A5. All digit pairs (e.g. D2 D’2) are attended to by P12.
Appendix J Appendix: N-Digit Addition
@@
Appendix K Appendix: Mixed Model Initialization
@@
Appendix L Appendix: Seed Sensitivity
An analysis of the sensitivity of 48 models to the initial seed was performed (This analysis excluded one model - the inaccurate 1-layer addition model that reproduces the Quirke and Barez [2024] paper results). Fig. 15 shows the results. We conclude:
  • The Addition models are the most stable - that is they are not sensitive to the seed value.
  • The other categories (Subtraction, Mixed, Mixed+Init, and Mixed+Reset) show relatively low to moderate sensitivity.
  • The higher average loss for Subtraction models show that the models find it harder to learn Subtraction is isolation.
  • The higher average loss for Mixed+Reset models show that the this type of intervention during training makes it harder for the models to learn
Figure 15: A visualization of the range of training losses across 48 models grouped by the different model categories.
Appendix M Appendix: N-Digit Subtraction
@@
Appendix N Appendix: Addition Hypothesis 3
  • Ablating the nodes one by one shows which answer digit(s) are reliant on each node (Ref Tab. 7). Most interestingly, ablating P10.L0.H1 impacts the answer digits A5, A4, A3, A2 (but not A1 and A0). This node is used in the calculation of A5, A4, A3, A2 in prediction steps 11, 12, 13 and 14. These relationships are constraints that are all obeyed by the pseudo-code.
  • The pseudo-code has 4 instances where $ST_{n}$ is calculated using TriCase. PCA of the corresponding nodes (P8.L0.H1, P9.L0.H1, P11.L0.H2 and P14.L0.H1) shows tri-state output for the specified $D_{n}$. (see Figure 7).
  • Activation patching (aka interchange intervention) experiments at attention head level confirmed some aspects of the calculations.
  • The pseudo code includes calculations like ST1 which it says is calculated in P9.L0.H1 and P9.L0.MLP. Ablation tells us both nodes are necessary. For the attention head we use the PCA results for insights. We didn’t implement a similar investigative tool for the MLP layer, so in the pseudo-code we attribute the calculation of ST1 to both nodes.
  • For P10.L0.H1, the attention head PCA could represent either a bi-state or tri-state output. The MLP layer at P10.L0.MLP could map the attention head output to either a bi-state or tri-state. We cannot see which. The pseudo-code shows a tri-state calculation at P10.L0.MLP, but with small alterations the pseudo-code would work with a bi-state output.
  • For P15.L0.H1 the attention head PCA could represent either a bi-state or tri-state output. The pseudo-code shows a bi-state calculation SC0 at P15.L0.H1, but with small alterations the pseudo-code would work with a tri-state output.
  • The calculation of ST2 in P14.L0.H1 is a interesting case. The model needs ST2 for A2 accuracy. The model could simply reuse the accurate ST2 value calculated in P10. Activation patching shows that it does not. Instead the P14 attention heads calculate ST1 from D1 and D’1 directly, and only relies on the P10.D1.ST2 value in the case where ST2 != ST1. That is, the calculation is “use P14.ST1 value else use ST2 value”. This aligns with the model learning the P10.ST1 calculation early in training (for 90% accuracy) and later learning that P10.ST2 contains additional information it can use to get to >99.999% accuracy.